Opened 9 years ago

Closed 3 years ago

#346 closed flight processing (fixed)

FW10/01, 187/2010, Dennys wood

Reported by: ella Owned by: knpa
Priority: alpha 4 low Milestone: 2010 data processing completion
Component: Archiving Keywords:
Cc: Other processors: ella

Description (last modified by emca)

Data location: ~arsf/arsf_data/2010/flight_data/uk/FW10_01-2010_187_Dennys_wood

Data arrived from ARSF via network transfer on:

Scientific objective: Do not have the ARSF Application pdf, but have requested this off Gary.

Priority: a4m

PI: Ross Hill

Sensors:

  • Eagle (24/08/2010) redelivered (13/12/2010)
  • Hawk (20/12/2010)
  • Leica LIDAR (14/09/2010)
  • Photography (18/08/2010)

Change History (53)

comment:1 Changed 9 years ago by jb362

Applanix processing complete and copied to workspace

comment:2 Changed 9 years ago by jb362

IPAS processing complete and copied to workspace

comment:3 Changed 9 years ago by mggr

  • Description modified (diff)
  • Priority changed from alpha 5 to alpha 4 low

Updated priority and PI per Gary's response. Still needs adding to processing priority list.

comment:4 Changed 9 years ago by benj

  • Description modified (diff)

comment:5 Changed 9 years ago by ella

Starting hyperspectral processing.

comment:6 Changed 9 years ago by emca

  • Description modified (diff)

comment:7 Changed 9 years ago by emca

  • Other processors set to emca
  • Owner set to emca
  • Status changed from new to assigned

comment:8 Changed 9 years ago by emca

Photography data ready for delivery check at
~airborne/workspace/FW10_01-2010_187_Dennys_wood/delivery/20100809

Removed photographs 1-12 as they did not cover the project area.

comment:9 Changed 9 years ago by jb362

Starting camera delivery check

comment:10 Changed 9 years ago by jb362

Camera delivery looks fine

comment:11 Changed 9 years ago by ella

Hyperspectral data ready for delivery check at ~arsf/workspace/FW10_01-2010_187_Dennys_wood/delivery/20100810

comment:12 Changed 9 years ago by knpa

Begun hyperspectral delivery check.

comment:13 Changed 9 years ago by ella

Flight line Eagle sct Hawk sct
01 0.20 0.60
02 -0.05 -0.08
03 0.01 0.03
04 -0.02 0.03
05 0.03 0.02
06 -0.05 -0.02
07 0.04 0.03
08 -0.02 -0.05
09 0.02 0.04
10 -0.02 -0.01
11 0.00 0.05
12 -0.02 -0.02
13 0.03 0.07
14 0.00 -0.01
15 0.04 0.02
16 0.00 0.00
17 1.00 1.00
18 1.00 1.40
19 0.05 0.01
20 0.05 0.01

comment:14 Changed 9 years ago by knpa

Checked. All looks good apart from the following:

  1. logsheet should be called FW10_01-2010_187_Dennys_wood.pdf
  2. No vectors. Check against vectors then:
    • Add screenshots of eagle/hawk mosaics with vectors
    • Update data quality section in config.
    • Update config to point to vectors to get vector images in readme
  3. Need to fix extra space by editing tex file:

Quality and Issues with data
"...we processed Level 3 imagery with a Digital Elevation Model . Screenshots of the Level 3..."

  1. Line 1 - lots of dropped lines at start - need to mention in Readme

Not checked Hawk due to waiting on bandset issue

comment:15 Changed 9 years ago by ella

  • Other processors changed from emca to ella

comment:16 Changed 9 years ago by ella

Everything mentioned in the above comment has now been sorted.

comment:17 Changed 9 years ago by mark1

  • Description modified (diff)

Photography dispatched to PI on 18/08/2010

comment:18 Changed 9 years ago by ella

  • Description modified (diff)

Eagle data dispatched to PI on 24th August 2010.

The hawk level 1 files are located in ~arsf/workspace/FW10_01-2010_187_Dennys_wood/delivery/hawk_20100823/lev1/

And the hawk jpg files and the mosaic with vector overlay are located in ~arsf/workspace/FW10_01-2010_187_Dennys_wood/delivery/hawk_20100823/screenshots/

comment:19 Changed 8 years ago by emca

LiDAR data ready for delivery check at
~airborne/workspace/FW10_01-2010_187_Dennys_wood/delivery/2010901

comment:20 Changed 8 years ago by jb362

Delivery checking LiDAR

comment:21 Changed 8 years ago by jb362

Delivery checked LiDAR. Every thing looks fine.

comment:22 Changed 8 years ago by emca

  • Description modified (diff)

LiDAR data sent to Ross Hill 14/10/2010

comment:23 Changed 8 years ago by emca

  • Description modified (diff)

Correction: LiDAR data sent to Ross Hill 14/09/2010

comment:24 Changed 8 years ago by knpa

  • Description modified (diff)

This project is in the repository and can be r-synced.

comment:25 Changed 8 years ago by emca

Corrected LAS1.3 files sent to R Hill 21/10/2010

(incorrect header on original delivered files due to an error in processing software)

comment:26 Changed 8 years ago by mark1

Hawk and Eagle time offsets appear to be incorrect, including the Eagle data previously dispatched. When the Hawk data is reprocessed also reprocess the Eagle and re-deliver.

comment:27 Changed 8 years ago by mark1

Update to hawk sct offsets:

line 1: 0.58
line 17: 1.25

comment:28 Changed 8 years ago by mark1

Have split the hawk lines 17 and 18 using azspec so as not to need -bend to process in azgcorr.

comment:29 Changed 8 years ago by mark1

Update to eagle sct offsets:

line 1: 0.35
line 18: 1.1

Have split the eagle lines 17 and 18 using azspec so as not to need -bend to process them in azgcorr

comment:30 Changed 8 years ago by emca

Have delivery checked the Hawk data - no issues.
Ready to deliver.

comment:31 Changed 8 years ago by emca

Have also delivery checked the Eagle data, that is good too :)

comment:32 Changed 8 years ago by mark1

  • Description modified (diff)

Hawk data has been dispatched to PI, as well as FW LiDAR extractions from the new extracting tool and Eagle re-delivery.

comment:33 Changed 8 years ago by emca

Hawk data has been checked and is ready for re-delivery

comment:34 Changed 8 years ago by emca

Hawk data re-delivered to PI 20/12/2010

comment:35 Changed 8 years ago by emca

  • Description modified (diff)

comment:36 Changed 8 years ago by mark1

project rsynced 02/02/2011. proc_laser not rsynced since LAS files included in delivery.

comment:37 Changed 8 years ago by emca

  • Component changed from Processing: general to Archiving

comment:38 Changed 7 years ago by knpa

  • Owner changed from emca to knpa
  • Status changed from assigned to accepted

Reprocessing hyperspectral with APL.

comment:39 Changed 7 years ago by knpa

Running stage 2.

comment:40 Changed 7 years ago by knpa

Re-running stage 2 (E+H) as apl chain has been updated to generate additional masked files.

comment:41 Changed 7 years ago by knpa

Re-running Hawk again as Mark made more changes in relation to bad pixel file.

Lines 17, 18a and 18b (both sensors) needed to go through again anyway due to mistakes with the config and issues with the split line.

comment:42 Changed 7 years ago by knpa

No longer need to split line 18 due to apl software handling the bends better.

New SCT values:

e17: 1.05
e18: 1.12 h18: 1.38

comment:43 Changed 7 years ago by knpa

More SCTs needed correct.

New values:

e08: -0.02
h17: 1.27

comment:44 Changed 7 years ago by knpa

  • Milestone set to 2010 data processing completion

Beginning archiving.

comment:45 Changed 7 years ago by knpa

Made available to NEODC 09/07/11

comment:46 Changed 7 years ago by knpa

*2012 I mean obviously

comment:47 Changed 7 years ago by adbe

Hyperspectral reprocessing delivery check:

  • gdal_translate command on page 4 looks incomplete
  • Should the bandlist for the example commands be ALL
  • Screenshots are at the end of the ReadME (remove [htb])
  • Mention dropped scans in the ReadMe

comment:48 Changed 7 years ago by knpa

Fixed above issues apart from mentioning dropped scans (of which there are none regular)

Preparing to deliver to Ross Hill.

comment:49 Changed 7 years ago by knpa

Delivered re-processed hyperspectral to Ross Hill in Bournemouth (24/07/12)

comment:50 Changed 6 years ago by knpa

Old deliveries ingested by NEODC.

Moved to archive disk.

comment:51 Changed 6 years ago by knpa

Deleted workspace version (original processing)

Leaving the ticket open until reprocessing is archived.

K

comment:52 Changed 3 years ago by dac

  • Status changed from accepted to new

Archiving

Data available from NEODC (http://browse.ceda.ac.uk/browse/neodc/arsf/2010/FW10_01/FW10_01-2010_187_Dennys_wood) - marking as closed.

comment:53 Changed 3 years ago by dac

  • Resolution set to fixed
  • Status changed from new to closed
Note: See TracTickets for help on using tickets.